Stock in Canacorp will pay a dividend of $1.23 at the end of one period, $2.45 at the end of two periods, and then dividends will grow at a constant rate of 6.25% per year indefinitely. If the required return is 11% we can value Canacorp stock by finding P2 using D3, then finding P0 = D1/(1.11) + D2/(1.11)2 + P2/(1.11)2. In this formula it almost appears as if we are ignoring all dividends from year three on. Discuss.
Answer #1

We are not ignoring. P2 includes all the dividends paid from year 3 onwards

D1 = 1.23

D2 = 2.45

D3 = 2.45 (1 + 6.25%) = 2.603125

P2 = D3 / required rate - growth rate

P2 = 2.603125 / 0.11 - 0.0625

P2 = 2.603125 / 0.0475

P2 = 54.802632

Current price = 1.23 / (1 + 0.11)1 + 2.45 / (1 + 0.11)2 + 54.802632 / (1 + 0.11)2

Current price = $47.58